OpenStudy (anonymous):

Solve the inequality: -7x + 4 ≤ 32

OpenStudy (anonymous):

-7x+4<=32 0<=7x+28 0<=x+4 i.e x>=-4

OpenStudy (anonymous):

awesome. thankyou